# Collaborative-Online-Learning-Enabled Distributionally Robust Motion Control for Multi-Robot Systems

### v1

@@ 1 Introduction
A plethora of methods have been proposed for collision avoidance motion control in robotic systems. We categorize and discuss these methods separately based on their application environments and target systems. Based on application environments, the methods can be classified into those designed for deterministic environments and those for uncertain environments. Collision avoidance in deterministic environments is a widely studied problem, with numerous well-established algorithms having been proposed, such as the penalty methods Hermans et al. (2021), control barrier function methods Dai et al. (2023), virtual holonomic constraints methods Otsason and Maggiore (2023), and among many others. Although the assumption of a deterministic environment simplifies the collision avoidance problem, it can incur safety issues in reality because prediction errors for obstacle positions are inevitable. Optimistically assuming perfect knowledge of future obstacle states may lead to unsafe trajectories when actual obstacle motions deviate from their predicted paths. Uncertain environments are more aligned with real-world scenarios, and our research is focused on this setting. In recent years, a large body of work on safe learning-based control has also been developed to address environmental uncertainty. For instance, Fisac et al. (2018) proposed a general framework that quantifies uncertainty from learning and ensures safety by augmenting the control space with safety margins. Yuan and Zhu (2024) developed a lightweight distributed Gaussian Process (GP) regression method for online multi-robot learning with limited bandwidth and computational power. Yuan et al. (2024) proposed a federated reinforcement learning approach to achieve zero-shot generalization in robot motion planning, offering theoretical guarantees on safety and performance across distributed agents. Beckers et al. (2021) proposed an online GP-based formation control law that learns unknown dynamics during runtime and guarantees bounded formation errors with high probability. Berkenkamp et al. (2016) developed a GP-based safe learning framework to estimate regions of attraction in uncertain nonlinear systems, enabling safe exploration with high-probability stability guarantees. For collision avoidance motion control in uncertain environments, chance-constrained methods are the most popular, ascribing to their capability to directly curb the chance of collision Hakobyan et al. (2019); Sharma et al. (2020). However, the safety probability guarantee of all the aforementioned chance-constrained methods heavily relies on the precise knowledge of distribution that is rarely available. To enhance the robustness against distribution errors, a novel control paradigm based on Distributionally Robust Optimization (DRO), called Distributionally Robust Model Predictive Control (DRMPC) was firstly proposed in Li et al. (2021) and Li et al. (2023a). Furthermore, a learning based data-driven MPC by using Conditional Value-at-Risk (CVaR) was proposed in Ning and You (2021). Building upon DRMPC, a novel motion control framework, referred to as Distributionally Robust Motion Control (DRMC) has attracted widespread attention, where the motion control is accomplished by solving a DRO problem at each control time. For the DRO problem, chance constraints should hold for all distributions within a set, called the ambiguity set Van Parys et al. (2015). Hakobyan and Yang (2021) utilized CVaR to measure safety risk and constructed a distance-based ambiguity set comprising all distributions within a Wasserstein ball centered at the empirical distribution. Building upon the work above, GP regression was employed to learn the motion distribution from real-time data and the estimated unimodal Gaussian distribution serves as the center of the Wasserstein ball Hakobyan and Yang (2022). Note that all of the methods above require solving a non-convex optimization problem at each control time, which is computationally intensive and unsuitable for real-time operation. Fortunately, the latest study Safaoui and Summers (2023) integrated the separating hyperplane technique with DRMC to convexify collision avoidance constraints, which enhances computational efficiency at the expense of control performance. The above literature review indicates that existing DRMC methods either entail substantial computational burdens or sacrifice a significant amount of control performance to enhance computational efficiency. DRMC that ensures safety while flexibly balancing computation time and control performance is currently lacking in academia. Moreover, all the aforementioned DRMC methods inherently impose restrictive assumptions, namely that the nominal trajectories of obstacles are the priori knowledge which is challenging to satisfy in practice.
@@ 2.2 Collision avoidance constraints
If constraints (8) and (9) are satisfied concurrently, constraint (7) is also satisfied. Subsequently, due to the uncertainty of the obstacle position $\varphi_{\ell}(k|t)$, enforcing the hard constraint (9) over its entire support set might be too conservative. Therefore, we denote the left-hand side of (9) as $\mathcal{L}_{\mathcal{H}_{i,\ell,t,k}^{obs}}(\varphi_{\ell}(k|t)\oplus\mathcal{O}_{\ell})$ and consider the following Distributionally Robust CVaR (DR-CVaR) constraint.
@@
where $\alpha^{u}$ is the used-defined confidence level and $\widetilde{\mathbb{D}}_{\ell,t,k}$ is the ambiguity set for $\varphi_{\ell}(k|t)$. If $\widetilde{\mathbb{D}}_{\ell,t,k}$ encapsulates the underlying distribution of $\varphi_{\ell}(k|t)$ and constraints (8) and (10) are simultaneously satisfied, constraint (7) is satisfied with a probability of at least $\alpha^{u}$.
@@ 2.3 Distributionally robust optimization for motion control
$\begin{split}&l_{i}(x_{i}(k|t),u_{i}(k|t),t)=\\ &\|x_{i}(k|t)-r_{i}^{x}(t+k)\|_{Q}+\|u_{i}(k|t)-r_{i}^{u}(t+k)\|_{R}\\ &q_{i}(x_{i}(K|t),t)=\|x_{i}(K|t)-r_{i}^{x}(t+K)\|_{P}\end{split}$ (12)
@@ 3.1 Background on DPMM
Our introduction to the DPMM is organized into two parts. The first part provides a detailed explanation of the classical DPMM theory. The second part presents the online variational inference algorithm that enables practical implementation in an online setting, along with its mathematical derivation.
@@ 3.1.1 DPMM theory
where $\mathrm{DP}$ represents the Dirichlet process (DP) with the concentration parameter $\alpha$ and the base distribution $H$ on the space of parameter $\phi$. $G$ is a discrete probability measure on the same space. $F(\phi)$ is the parametric distribution family with $\phi$ being its parameters and the data point $\omega$ is sampled from the distribution $F(\phi)$. Drawing $G$ from $\mathrm{DP}(\alpha,H)$ admits the following stick-breaking procedure Sethuraman (1994).
@@
Given the weighted sum expression of discrete distribution $G$ as in (15), drawing a sample $\phi_{n}$ from $G$ can be viewed as picking a mass point from $\phi_{n}$ whose probability is $\gamma_{n}$. Let the latent variable $z$ denote the selected mass point. Since each component $\phi_{n}$ is chosen with probability $\gamma_{n}$, the variable $z$ follows a multinomial distribution over the weight vector $\bm{\gamma}$, denoted as follows.
@@ 3.1.2 Online variational inference algorithm of DPMM
where $KL(\cdot|\cdot)$ denotes the Kullback–Leibler divergence, $K$ is the number of mixture components, $N$ the total data volume, $T$ the number of processed observations, and $n_{s}$ the size of clump $s$. The term $S_{sk}$ measures the statistical affinity between clump $s$ and component $k$. Based on this clump-level objective (20), the algorithm recursively partitions the data to identify optimal clump structures. The full procedure is detailed in Algorithm 1.
@@ 4.1 Collaborative online learning method
Note that we choose $F(\cdot)$ as the Gaussian distribution family, and thus at time $t$, the learning result for the obstacle motion $\omega_{\ell,t}$ includes the number of the mixture components $m_{\ell,t}$ and for each component $i$: (1) the number of learning data $N$, (2) the estimated weight $\hat{\gamma}$, and (3) the estimated mean $\hat{\mu}$ and covariance $\hat{\Sigma}$ which are essentially computed by (24).
$\begin{split}&\hat{\mu}=\frac{1}{N}\sum\nolimits_{j=1}^{N}\hat{\omega}_{\ell}^{i(j)}\\ &\hat{\Sigma}=\frac{1}{N}\sum\nolimits_{j=1}^{N}(\hat{\omega}_{\ell}^{i(j)}-\hat{\mu})(\hat{\omega}_{\ell}^{i(j)}-\hat{\mu})^{T}\end{split}$ (24)
where $\hat{\omega}_{\ell}^{i(j)}$ is the learning data from component $i$.
@@ 4.2 Ambiguity set construction and propagation
Based on the results of COOL, we develop an online-updated ambiguity set for $\omega_{\ell,t}$. The definition of this ambiguity set is first stated explicitly, followed by a detailed explanation.
@@
where $\bm{\gamma_{\ell,t}}=(\gamma_{\ell,t}^{1},...,\gamma_{\ell,t}^{m_{\ell,t}})$ and $\bm{\hat{\gamma}_{\ell,t}}=(\hat{\gamma}_{\ell,t}^{1},...,\hat{\gamma}_{\ell,t}^{m_{\ell,t}})$. $\mathbb{D}_{\ell,t}$ in (25c) is the ambiguity set devised as a weighted Minkowski sum of $m_{\ell,t}$ basic ambiguity sets $\mathbb{D}$. Intuitively, $\mathbb{D}_{\ell,t}$ incorporates two layers of robustness: the first accounts for the uncertainty in the estimated weights, modeled by the set $\Delta_{\ell,t}$, and the second addresses the uncertainty in the estimated moments, captured by the structure of the basic ambiguity set $\mathbb{D}$. Specifically, $\Delta_{\ell,t}$ in (25d) is the confidence region of the weight $\bm{\gamma_{\ell,t}}$, where $\theta_{\ell,t}$ is a parameter indicating the total variation distance. For the definition of basic ambiguity set $\mathbb{D}$ in (25e), the second constraint specifies that the first-order moment of any distribution in $\mathbb{D}$ belongs to a confidence region scaled by the parameter $\beta$. Similarly, the third constraint ensures that the second-order moment lies within a confidence region scaled by the parameter $\varepsilon$. $\hat{\mu}$ and $\hat{\Sigma}$ are estimated first-order and second-order moments, respectively. $\mathcal{M}(\mathbb{W}_{\omega})$ is the set of positive Borel measures on $\mathbb{W}_{\omega}$.
@@
(Finite-sample guarantee) Suppose Assumption 7 holds, given the confidence level $\chi_{\ell}$ and $\alpha_{\ell}^{i}$, let $\hat{\alpha}_{\ell,t}^{i}=1-\sqrt{\alpha_{\ell}^{i}}$, $\overline{R}_{\ell,t}^{i}=\left(1-\left(\left(\hat{R}_{\ell,t}^{i}\right)^{2}+2\right)\left(2+\sqrt{2\ln\left(4/\hat{\alpha}_{\ell}^{i}\right)}\right)\right$.
$\left.\Big{/}\sqrt{N_{\ell,t}^{i}}\right)^{-\frac{1}{2}}\hat{R}_{\ell,t}^{i}$ and $\hat{R}_{\ell,t}^{i}=\sup_{\omega\in\mathbb{W}_{\omega}}\|(\hat{\Sigma}_{\ell,t}^{i})^{-\frac{1}{2}}(\omega-\hat{\mu}_{\ell,t}^{i})\|_{2}$, while the parameters $\theta_{\ell,t}$, $\beta$ and $\varepsilon$ of the ambiguity set $\mathbb{D}_{\ell,t}$ are defined as follows (the subscripts of $\hat{\alpha}_{\ell}^{i}$, $\overline{R}_{\ell,t}^{i}$, $\hat{R}_{\ell,t}^{i}$, and part of $N_{\ell,t}^{i}$ are omitted)
@@
$\varepsilon_{\ell,t,k}^{j}=\beta_{\ell,t,k}^{j}+\max_{i=1,...,m_{\ell,t}}\{\varepsilon-\beta\}$ (41g)

### v2

@@ 1 Introduction
A plethora of methods have been proposed for collision avoidance motion control in robotic systems. We categorize and discuss these methods separately based on their application environments and target systems. Based on application environments, the methods can be classified into those designed for deterministic environments and those for uncertain environments. Collision avoidance in deterministic environments is a widely studied problem, with numerous well-established algorithms having been proposed, such as the penalty methods Hermans et al. (2021), control barrier function methods Dai et al. (2023), virtual holonomic constraints methods Otsason and Maggiore (2023), and among many others. Although the assumption of a deterministic environment simplifies the collision avoidance problem, it can incur safety issues in reality because prediction errors for obstacle positions are inevitable. Optimistically assuming perfect knowledge of future obstacle states may lead to unsafe trajectories when actual obstacle motions deviate from their predicted paths. Uncertain environments are more aligned with real-world scenarios, and our research is focused on this setting. In recent years, a large body of work on safe learning-based control has also been developed to address environmental uncertainty. For instance, Fisac et al. (2018) proposed a general framework that quantifies uncertainty from learning and ensures safety by augmenting the control space with safety margins. Yuan and Zhu (2024) developed a lightweight distributed Gaussian Process (GP) regression method for online multi-robot learning with limited bandwidth and computational power. Yuan and Zhu (2025) proposed a distributed safe learning and planning framework for multi-robot systems, which leverages GP to model external disturbance dynamics and derive the sufficient conditions to guarantee the safety. Beckers et al. (2021) proposed an online GP-based formation control law that learns unknown dynamics during runtime and guarantees bounded formation errors with high probability. Berkenkamp et al. (2016) developed a GP-based safe learning framework to estimate regions of attraction in uncertain nonlinear systems, enabling safe exploration with high-probability stability guarantees. For collision avoidance motion control in uncertain environments, chance-constrained methods are the most popular, ascribing to their capability to directly curb the chance of collision Hakobyan et al. (2019); Sharma et al. (2020). However, the safety probability guarantee of all the aforementioned chance-constrained methods heavily relies on the precise knowledge of distribution that is rarely available. To enhance the robustness against distribution errors, a novel control paradigm based on Distributionally Robust Optimization (DRO), called Distributionally Robust Model Predictive Control (DRMPC) was firstly proposed in Li et al. (2021) and Li et al. (2023a). Furthermore, a learning based data-driven MPC by using Conditional Value-at-Risk (CVaR) was proposed in Ning and You (2021). Building upon DRMPC, a novel motion control framework, referred to as Distributionally Robust Motion Control (DRMC) has attracted widespread attention, where the motion control is accomplished by solving a DRO problem at each control time. For the DRO problem, chance constraints should hold for all distributions within a set, called the ambiguity set Van Parys et al. (2015). Hakobyan and Yang (2021) utilized CVaR to measure safety risk and constructed a distance-based ambiguity set comprising all distributions within a Wasserstein ball centered at the empirical distribution. Building upon the work above, GP regression was employed to learn the motion distribution from real-time data and the estimated unimodal Gaussian distribution serves as the center of the Wasserstein ball Hakobyan and Yang (2022). Note that all of the methods above require solving a non-convex optimization problem at each control time, which is computationally intensive and unsuitable for real-time operation. Fortunately, the latest study Safaoui and Summers (2023) integrated the separating hyperplane technique with DRMC to convexify collision avoidance constraints, which enhances computational efficiency at the expense of control performance. The above literature review indicates that existing DRMC methods either entail substantial computational burdens or sacrifice a significant amount of control performance to enhance computational efficiency. DRMC that ensures safety while flexibly balancing computation time and control performance is currently lacking in academia. Moreover, all the aforementioned DRMC methods inherently impose restrictive assumptions, namely that the nominal trajectories of obstacles are the priori knowledge which is challenging to satisfy in practice.
@@ 2.2 Collision avoidance constraints
If constraints (8) and (9) are satisfied concurrently, constraint (7) is also satisfied. We note that if the position of the obstacle $\ell$ at the predicted time $k|t$ is known, $h_{i,\ell,t,k}^{obs}$ can be directly computed as follows.
$\begin{split}h_{i,\ell,t,k}^{obs}=\frac{\varphi_{\ell}(k|t)-y_{i}(k|t)}{\|\varphi_{\ell}(k|t)-y_{i}(k|t)\|_{2}}\end{split}$ (10)
However, due to the uncertainty of $\varphi_{\ell}(k|t)$, $h_{i,\ell,t,k}^{obs}$ cannot be directly computed using (10). Instead, $\varphi_{\ell}(k|t)$ or its distribution needs to be learned or estimated from data. Therefore, we denote the left-hand side of (9) as $\mathcal{L}_{\mathcal{H}_{i,\ell,t,k}^{obs}}(\varphi_{\ell}(k|t)\oplus\mathcal{O}_{\ell})$ and consider the following Distributionally Robust CVaR (DR-CVaR) constraint.
@@
where $\alpha^{u}$ is the used-defined confidence level and $\widetilde{\mathbb{D}}_{\ell,t,k}$ is the ambiguity set for $\varphi_{\ell}(k|t)$. Additionally, ${\rm CVaR}_{\alpha}^{\upsilon}(X)=\min_{z\in\mathbb{R}}\mathbb{E}^{\upsilon}\left[z+(X-z)^{+}/(1-\alpha)\right]$ measures the conditional expectation of the loss within the $(1-\alpha)$ worst-case quantile. If $\widetilde{\mathbb{D}}_{\ell,t,k}$ encapsulates the underlying distribution of $\varphi_{\ell}(k|t)$ and constraints (8) and (11) are simultaneously satisfied, constraint (7) is satisfied with a probability of at least $\alpha^{u}$.
It is evident that when constraints (8) and (11) are simultaneously satisfied, the distributionally robust collision avoidance requirement between agent $i$ and obstacle $\ell$ is guaranteed. However, since obstacles are not controllable, constraint (11) cannot be directly enforced during the motion control problem. Instead, constraint (11) is taken into account in the computation of the separating hyperplane $\mathcal{H}_{i,\ell,t,k}^{obs}$, while constraint (8) is explicitly enforced in the motion control problem of agent $i$. Through this decoupled treatment, the distributionally robust agent–obstacle collision avoidance requirement is effectively ensured.
@@ 2.3 Distributionally robust optimization for motion control
$\begin{split}&l_{i}(x_{i}(k|t),u_{i}(k|t))=\\ &\|x_{i}(k|t)-r_{i}^{x}(t+k)\|_{Q}+\|u_{i}(k|t)-r_{i}^{u}(t+k)\|_{R}\\ &q_{i}(x_{i}(K|t))=\|x_{i}(K|t)-r_{i}^{x}(t+K)\|_{P}\end{split}$ (13)
@@ 3.1 Background on DPMM
In this paper, the DPMM is used to model the stochastic motion of obstacles by learning the probability distribution of their inter-sample displacements. In this subsection, our introduction to the DPMM is organized into two parts. The first part provides a detailed explanation of the classical DPMM theory. The second part presents the online variational inference algorithm that enables practical implementation in an online setting, along with its mathematical derivation.
@@ 3.1.1 DPMM theory
where $\mathrm{DP}$ represents the Dirichlet process (DP) with the concentration parameter $\alpha$ and the base distribution $H$ on the space of parameter $\phi$. $G$ is a discrete probability measure on the same space. $F(\phi)$ is the parametric distribution family with $\phi$ being its parameters and the data point $\omega$ is sampled from the distribution $F(\phi)$. Here, the data point $\omega$ corresponds to the observed displacement of an obstacle between two consecutive time steps, capturing its random motion over one sampling interval. Drawing $G$ from $\mathrm{DP}(\alpha,H)$ admits the following stick-breaking procedure Sethuraman (1994).
@@
Given the weighted sum expression of discrete distribution $G$ as in (16), drawing a sample $\phi_{n}$ from $G$ can be viewed as picking a component whose probability is $\gamma_{n}$. Let the latent variable $z$ denote the selected mass point. In this paper, each component corresponds to a distinct obstacle motion mode, which represents one possible stochastic behavior of the obstacle over a single sampling interval. Since each component $\phi_{n}$ is chosen with probability $\gamma_{n}$, the variable $z$ follows a multinomial distribution over the weight vector $\bm{\gamma}$, denoted as follows.
@@ 3.1.2 Online variational inference algorithm of DPMM
where $KL(\cdot|\cdot)$ denotes the Kullback–Leibler divergence, $K$ is the number of mixture components, $N$ the total data volume, $T$ the number of processed observations, and $n_{s}$ the size of clump $s$. The term $S_{sk}$ measures the statistical affinity between clump $s$ and component $k$. Based on this clump-level objective (21), the algorithm recursively partitions the data to identify optimal clump structures. It is important to note that clumps are algorithmic structures introduced for efficient online inference and information aggregation, and do not correspond to any physical quantity of obstacle motion. The full procedure is detailed in Algorithm 1.
@@ 4.1 Collaborative online learning method
Note that we choose $F(\cdot)$ as the Gaussian distribution family, and thus at time $t$, the learning result for the obstacle motion $\omega_{\ell,t}$ includes the number of the mixture components $m_{\ell,t}$ and for each component $i$: (1) the number of learning data $N_{\ell,t}^{i}$, (2) the estimated weight $\hat{\gamma}_{\ell,t}^{i}$, and (3) the estimated mean $\hat{\mu}_{\ell,t}^{i}$ and covariance $\hat{\Sigma}_{\ell,t}^{i}$ which are essentially computed by (25).
$\begin{split}&\hat{\mu}_{\ell,t}^{i}=\frac{1}{N_{\ell,t}^{i}}\sum\nolimits_{j=1}^{N_{\ell,t}^{i}}\hat{\omega}_{\ell,t}^{i(j)}\\ &\hat{\Sigma}_{\ell,t}^{i}=\frac{1}{N_{\ell,t}^{i}}\sum\nolimits_{j=1}^{N_{\ell,t}^{i}}(\hat{\omega}_{\ell,t}^{i(j)}-\hat{\mu}_{\ell,t}^{i})(\hat{\omega}_{\ell,t}^{i(j)}-\hat{\mu}_{\ell,t}^{i})^{T}\end{split}$ (25)
where $\hat{\omega}_{\ell,t}^{i(j)}$ is the learning data from component $i$.
@@ 4.2 Ambiguity set construction and propagation
Based on the results of COOL, we develop an online-updated ambiguity set for $\omega_{\ell,t}$. This ambiguity set is designed to characterize the uncertainty in the probability distribution of the obstacle motion vector $\omega_{\ell,t}$ learned from data, and serves as the foundation for constructing distributionally robust collision avoidance constraints. The definition of this ambiguity set is first stated explicitly, followed by a detailed explanation.
@@
where $\bm{\gamma_{\ell,t}}=(\gamma_{\ell,t}^{1},...,\gamma_{\ell,t}^{m_{\ell,t}})$ and $\bm{\hat{\gamma}_{\ell,t}}=(\hat{\gamma}_{\ell,t}^{1},...,\hat{\gamma}_{\ell,t}^{m_{\ell,t}})$. $\mathbb{D}_{\ell,t}$ in (26c) is the ambiguity set devised as a weighted Minkowski sum of $m_{\ell,t}$ basic ambiguity sets $\mathbb{D}$. Intuitively, $\mathbb{D}_{\ell,t}$ incorporates two layers of robustness: the first accounts for the uncertainty in the estimated weights, modeled by the set $\Delta_{\ell,t}$, and the second addresses the uncertainty in the estimated moments, captured by the structure of the basic ambiguity set $\mathbb{D}_{\ell,t}^{i}$. Specifically, $\Delta_{\ell,t}$ in (26d) is the confidence region of the weight $\bm{\gamma_{\ell,t}}$, where $\theta_{\ell,t}$ is a parameter indicating the total variation distance. For the definition of basic ambiguity set $\mathbb{D}_{\ell,t}^{i}$ in (26e), the second constraint specifies that the first-order moment of any distribution in $\mathbb{D}_{\ell,t}^{i}$ belongs to a confidence region scaled by the parameter $\beta_{\ell,t}^{i}$. Similarly, the third constraint ensures that the second-order moment lies within a confidence region scaled by the parameter $\varepsilon_{\ell,t}^{i}$. $\hat{\mu}_{\ell,t}^{i}$ and $\hat{\Sigma}_{\ell,t}^{i}$ are estimated first-order and second-order moments, respectively. $\mathcal{M}(\mathbb{W}_{\omega})$ is the set of positive Borel measures on $\mathbb{W}_{\omega}$. From a physical perspective, the ambiguity set $\mathbb{D}_{\ell,t}$ represents a family of plausible probabilistic models for the obstacle displacement $\omega_{\ell,t}$, capturing the uncertainty in both the learned motion modes and their relative likelihoods.
@@
(Finite-sample guarantee) Suppose Assumption 7 holds, given the confidence level $\chi_{\ell}$ and $\alpha_{\ell}^{i}$, let $\hat{\alpha}_{\ell,t}^{i}=1-(\alpha_{\ell}^{i})^{\frac{1}{2}}$, $\overline{R}_{\ell,t}^{i}=(1-((\hat{R}_{\ell,t}^{i})^{2}+2)(2+(2\ln(4/\hat{\alpha}_{\ell}^{i}))^{\frac{1}{2}})/(N_{\ell,t}^{i})^{\frac{1}{2}})^{-\frac{1}{2}}\hat{R}_{\ell,t}^{i}$ and $\hat{R}_{\ell,t}^{i}=\sup_{\omega\in\mathbb{W}_{\omega}}\|(\hat{\Sigma}_{\ell,t}^{i})^{-\frac{1}{2}}(\omega-\hat{\mu}_{\ell,t}^{i})\|_{2}$, while the parameters $\theta_{\ell,t}$, $\beta$ and $\varepsilon$ of the ambiguity set $\mathbb{D}_{\ell,t}$ are defined as follows (the subscripts of $\hat{\alpha}_{\ell}^{i}$, $\overline{R}_{\ell,t}^{i}$, $\hat{R}_{\ell,t}^{i}$, and part of $N_{\ell,t}^{i}$ are omitted)
@@
$\varepsilon_{\ell,t,k}^{j}=\beta_{\ell,t,k}^{j}+\max_{i=1,...,m_{\ell,t}}\varepsilon_{\ell,t}^{i}$ (42g)